OpenStudy (anonymous):

x/2 = 5 x=10 How can you tell if lines are parallel, perpendicular, or neither?

OpenStudy (kymber):

y=mx+b, M is slope x/2=5, slope is 1/2 x = 10, slope is 1 If the lines are parallel, the slope will be the exact same. If they're perpendicular, then the slopes will be opposite reciprocals. These are neither. :]

OpenStudy (anonymous):

x/3=3 x=9 How can you tell if lines are parallel, perpendicular, or neither?

OpenStudy (kymber):

 These are neither